# {Titlepage}

### v1

@@
The discovery of a terrestrial-mass free-floating planet candidate in the light curve of the star TIC 107150013 observed by the Transiting Exoplanet Survey Satellite (TESS) has recently been announced. A short-duration ($\sim 0.5$ day), low-amplitude ($\sim 0.06$ mag) brightening in the TESS light curve was interpreted as a short-timescale gravitational microlensing event. However, the purported event occurred far from the Galactic center and the Galactic plane ($l\approx 239^{\circ}$, $b\approx-5^{\circ}$), on a relatively nearby ($\sim 3.2$ kpc) star, making the microlensing interpretation unlikely. Here, we report the archival photometric observations of TIC 107150013 collected by the Optical Gravitational Lensing Experiment (OGLE) from 2018 through 2020. The archival OGLE light curve reveals periodic variability indicative of starspots on the surface of the star. The presence of starspots indicates magnetic activity of the star, which may also manifest as stellar flares. We interpret the brightening of TIC 107150013 seen in the TESS data as the stellar flare. We present similar flaring stars detected in the archival OGLE data, mimicking short-timescale, low-amplitude microlensing events. Such stars may be a source of non-negligible false positive detections in the planned space-based microlensing surveys. Gravitational microlensing (672), Free-floating planets (549), Stellar activity (1580), Starspots (1572), Stellar flares (1603), Flare stars (540)
@@ 1 Introduction
Only one object, TIC 107150013, met all selection criteria and was considered to be a compelling microlensing event candidate. The brightening observed by TESS lasted about $\sim 0.5$ day and had a low amplitude ($\sim 0.06$ mag). The best-fit microlensing model fitted to the light curve yielded the Einstein timescale $t_{\rm E}=0.074\pm 0.002$ days, and the angular Einstein radius $\theta_{\rm E}=4.1^{+0.4}_{-0.5}\,\mu\mathrm{as}$, which are similar to those of other known free-floating planet candidates.
All known microlensing events due to free-floating planets have occurred toward the dense regions of the Galactic bulge, where the probability of finding microlensing events is highest. On the contrary, the star TIC 107150013 is located at the Galactic coordinates $l=239.1^{\circ}$, $b=-5.0^{\circ}$, at a distance of $3.19\pm 0.15$ kpc (Kunimoto et al., 2024). The probability of observing a microlensing event so far away from the Galactic center and the Galactic plane, on such a nearby star, is extremely low. Mróz et al. (2020) found that the microlensing event rate in the Galactic plane decreases exponentially with the increasing angular separation from the Galactic center $\Gamma\propto\exp(-|l|/l_{0})$, where $l_{0}=31.5^{+4.2}_{-3.7}$ deg. At $l=-120.9^{\circ}$, the average event rate is approximately $0.6\times 10^{-7}\,\mathrm{yr}^{-1}\,\mathrm{star}^{-1}$, and is $\sim 300$ times lower than that toward the Galactic bulge. The estimated number of microlensing events in one TESS sector is approximately $0.6\times 10^{-7}\times 1.3\times 10^{6}\times 27.4/365.25\approx 0.006$, assuming 100% detection efficiency. The expected number of microlensing events is even lower because Mróz et al. (2020) calculated the event rate averaged over the latitude range $-7^{\circ}<b<7^{\circ}$ and $\Gamma$ is known to be decreasing with the increasing distance from the Galactic plane, the analyzed TESS star is relatively nearby, and events due to free-floating planets are expected to be far less frequent than those by stars. That makes the interpretation that TIC 107150013 showed a short-duration microlensing event very unlikely.
@@ 2 Archival OGLE Observations of TIC 107150013
Even though the rotation period of the star is longer than the duration of TESS observations, a low-amplitude, long-timescale variability can be noticed in the pre-detrended TESS data, which are presented in Figure 2 of Kunimoto et al. (2024).
@@ 3 Discussion
We analyzed the light curves of 79 flaring stars published by Iwanek et al. (2019) and found two objects with flares similar to that seen in TIC 107150013 (Figure 2). Two such flaring episodes occurred for the left panel object over the last 14 years of the OGLE monitoring and just one for the right panel star. In both cases, an additional periodic variability due to starspots is present in the data. Therefore, a natural explanation of the brightening observed in the light curve of TIC 107150013 is a stellar flare, not a short-timescale, low-amplitude microlensing event.

### v2

@@
The discovery of a terrestrial-mass free-floating planet candidate in the light curve of the star TIC 107150013 observed by the Transiting Exoplanet Survey Satellite (TESS) has recently been announced. A short-duration ($\approx 0.5$ day), low-amplitude ($\approx 0.06$ mag) brightening in the TESS light curve was interpreted as a short-timescale gravitational microlensing event. However, the purported event occurred far from the Galactic center and the Galactic plane ($l\approx 239^{\circ}$, $b\approx-5^{\circ}$), on a relatively nearby ($\approx 3.2$ kpc) star, making the microlensing interpretation unlikely. Here, we report the archival photometric observations of TIC 107150013 collected by the Optical Gravitational Lensing Experiment (OGLE) from 2018 through 2020. The archival OGLE light curve reveals periodic variability indicative of starspots on the surface of the star. The presence of starspots indicates magnetic activity of the star, which may also manifest as stellar flares. We interpret the brightening of TIC 107150013 seen in the TESS data as the stellar flare. We present similar flaring stars detected in the archival OGLE data, mimicking short-timescale, low-amplitude microlensing events. Such stars may be a source of non-negligible false positive detections in the planned space-based microlensing surveys. Gravitational microlensing (672), Free-floating planets (549), Stellar activity (1580), Starspots (1572), Stellar flares (1603), Flare stars (540)
@@ 1 Introduction
Only one object, TIC 107150013 ($\mathrm{RA}=07^{\rm h}22^{\rm m}30\mbox{$.\!\!^{\mathrm{s}}$}81$, $\mathrm{Decl.}={-25}\mbox{${}^{\circ}$}{29}\mbox{${}^{\prime}$}{15}\mbox{$.\!\!^{\prime\prime}$}{7}$, J2000), met all selection criteria and was considered to be a compelling microlensing event candidate. The brightening observed by TESS lasted about $\approx 0.5$ day and had a low amplitude ($\approx 0.06$ mag). The best-fit microlensing model fitted to the light curve yielded the Einstein timescale $t_{\rm E}=0.074\pm 0.002$ days, and the angular Einstein radius $\theta_{\rm E}=4.1^{+0.4}_{-0.5}\,\mu\mathrm{as}$, which are similar to those of other known free-floating planet candidates.
All known microlensing events due to free-floating planets have occurred toward the dense regions of the Galactic bulge, where the probability of finding microlensing events is highest. On the contrary, the star TIC 107150013 is located at the Galactic coordinates $l=239.1^{\circ}$, $b=-5.0^{\circ}$, at a distance of $3.19\pm 0.15$ kpc (Kunimoto et al., 2024). The probability of observing a microlensing event so far away from the Galactic center and the Galactic plane, on such a nearby star, is extremely low. Mróz et al. (2020) found that the microlensing event rate in the Galactic plane decreases exponentially with the increasing angular separation from the Galactic center $\Gamma\propto\exp(-|l|/l_{0})$, where $l_{0}=31.5^{+4.2}_{-3.7}$ deg. At $l=-120.9^{\circ}$, the average event rate is approximately $0.6\times 10^{-7}\,\mathrm{yr}^{-1}\,\mathrm{star}^{-1}$, and is $\approx 300$ times lower than that toward the Galactic bulge. The estimated number of microlensing events in one TESS sector is approximately $0.6\times 10^{-7}\times 1.3\times 10^{6}\times 27.4/365.25\approx 0.006$, assuming 100% detection efficiency. The expected number of microlensing events is even lower because Mróz et al. (2020) calculated the event rate averaged over the latitude range $-7^{\circ}<b<7^{\circ}$ and $\Gamma$ is known to be decreasing with the increasing distance from the Galactic plane, the analyzed TESS star is relatively nearby, and events due to free-floating planets are expected to be far less frequent than those by stars. Indeed, a detailed calculation by Yang et al. (2024) found that only 0.0018 microlensing events due to free-floating planets are expected to be detected in TESS Sector 61. That makes the interpretation that TIC 107150013 showed a short-duration microlensing event very unlikely.
@@ 2 Archival OGLE Observations of TIC 107150013
Even though the rotation period of the star is longer than the duration of TESS observations, a low-amplitude, long-timescale variability can be noticed in the pre-detrended TESS data, which are presented in Figure 2 of Kunimoto et al. (2024). Moreover, variability with a period of 43–44 days can be clearly detected in the publicly available light curve of TIC 107150013 from the All-Sky Automated Survey for Supernovae (Shappee et al., 2014; Hart et al., 2023).
@@ 3 Discussion
We analyzed the light curves of 79 flaring stars published by Iwanek et al. (2019) and found two objects with flares similar to that seen in TIC 107150013 (Figure 2). Two such flaring episodes occurred for the left panel object over the last 14 years of the OGLE monitoring and just one for the right panel star. Taking into account diurnal and seasonal gaps in the data, the frequency of flares in these stars is $1\div 2$ yr-1. In both cases, an additional periodic variability due to starspots is present in the data. Moreover, a close inspection of light curve of TIC 107150013 seen in Figure 5 of Kunimoto et al. (2024) (especially improved reductions in the left panel) reveals a fine structure of the observed brightening, similar to flickering in stellar eruptive episodes, and unexpected in smooth light curves of microlensing events. Therefore, a natural explanation of the brightening observed in the light curve of TIC 107150013 is a stellar flare, not a short-timescale, low-amplitude microlensing event.
TIC 107150013 is a giant, and stellar flares are known to be relatively rare in such stars. In a study based on Kepler observations, Van Doorsselaere et al. (2017) found that about 3.18% of giants are flaring, an incidence rate similar to that observed in main-sequence stars. However, later studies by Yang and Liu (2019) and Oláh et al. (2021) revealed that in many cases, giant oscillations or pulsations were mistaken with flares by an automated algorithm of Van Doorsselaere et al. (2017). These refined analyses found that only $0.3\%$ giants observed by Kepler exhibit flares.
Moreover, Oláh et al. (2021) found that the flaring activity of giants is always connected to their rotational variability. Flares were observed in giants with the photometric index $S_{\rm ph}$ higher than 0.2%, whereas the typical inactive red giants have $S_{\rm ph}<0.1\%$ (Gaulme et al., 2020). (The photometric index is essentially the mean standard deviation of the time series data.) In the case of TIC 107150013, we measured $S_{\rm ph}=1.7\%$, which is typical for other known flaring giants (Oláh et al., 2021).
The sheer number of giants observed by TESS makes the stellar flare interpretation much more likely than microlensing. We queried the revised TESS Input Catalog (TIC, v8.2; Stassun et al. 2018, 2019; Paegert et al. 2021) to get radii and effective temperatures of $10^{5}$ randomly selected stars from TESS Sector 61. We found that about 38% are giants (using the same cut on effective temperature and radius as that adopted in Figure 11 of Stassun et al. 2019). Therefore, the expected number of flares on giant stars is on the order of $0.3\%\times 0.38\times 1.3\times 10^{6}\times 1\times 27.4/365.25\approx 100$, a factor of $10^{5}$ larger than the expected number of free-floating planets in the same data set. Most of these flaring objects were rejected by the algorithm of Kunimoto et al. (2024), except only one candidate.